I keep getting this question wrong, any help?

Supposing that you need to clamp a 5/16 thick steel plate with a 3/4 thick brass plate using a 1/4-20 bolt, what length bo
Answer #1

Thickness of Steel Plate, t_s =5/16"

Thickness of Brass plate, t_b =3/6"

Thickness of regulat 1/4"-20 nut, t_{nut} =7/32"

Thickness of washer for 1/4" - 20 bolt, t_{washer} =0.065"

The bolt should project upto a maximum of 3 threads, so since the bolt is of 20 tpi, 3 threads should have a length of 0.15 in

Total thickness , t_s+t_b+t_{nut}+t_{washer}+0.15=1.24625 \approx 1.25